A lake has still waters (no moving current). A boat on that lake can move at a speed of 7.5 km/h. If the boat is now moving on a stream that has flowing water, determine how fast the boat can move, relative to a tree on shore, when it travels (a) upstream and (b) downstream. Assume the stream has a speed of 3.2 km/h.
Answer #1

A)

Relative speed, V = u - v = 7.5 - 3.2 = 4.3 km/h

B)

Relative speed, V = u + v = 7.5 + 3.2 = 10.9 km/h

Comment in case any doubt please rate my answer...